OpenStudy (anonymous):

0.19494949494.....in a fraction???

OpenStudy (tkhunny):

It might be easier to write 1.9494949494... As a fraction. 1 + 94/99 = (99+94)/99 = 193/99 Done Okay, now how do we answer the question?

OpenStudy (perl):

you can multiply this by the period

myininaya (myininaya):

Example: This is my favorite way: Say we wanted to write .331331331331331... as a fraction. \[\text{ Let } x= .331331331... \\ \text{ so } 1000x=331.331331... \\ 1000x-x=331.331331...-.331331331 \\ 999x=331 \\ x=\frac{331}{999} \\ \]

OpenStudy (anonymous):

@perl are you saying 194/99?right

OpenStudy (perl):

not exactly

OpenStudy (perl):

i dont think you have to find the exact rational number the question says, is this a rational number It is a mathematical fact that repeating decimals are rational numbers.

OpenStudy (perl):

eventually repeating decimals*

OpenStudy (perl):

mathematical fact = theorem (which means we can prove it)

OpenStudy (perl):

I'll just finish teh argument if x = 0.19494949494.... then 10x = 1.949494... 1000x = 194.949494... 1000x - 10x = 193 990x = 193 x = 193/990

OpenStudy (perl):

ok i see my mistake earlier, there is a nice algorithm .19494949 = .1 + .0949494.... = 1/10 + 94/990

OpenStudy (perl):

so if you have an eventually repeating decimal cut out the part which is not repeating and add it the repeating decimal.

OpenStudy (perl):

lets do another example consider the decimal .13678678678... we see that '678' repeats .13678678678... = .13 + .00678678... = 13/100 + 678/99900